# Defining a Role-Centered Terminology for Physical Representations and Controls

### v1

@@ 10. Voodoo Dolls 1998 (Isidoro and Sclaroff, 1998) Spongy object $\times$ - - - - $\times$ - DATABLE
11. mediaBlocks 1998 (Ullmer and Ishii, 1999; Ullmer et al., 1998) mediaBlock - - $\times$ - - $\times$ - OPABLE
@@
12. musicBottles 1999 (Ishii et al., 1999, 2001a) Music $\times$ - - - - - $\times$ DATNIBLE
@@
13. Urp (Urban Planning Workbench) 1999 (Ben-Joseph et al., 2001; Ishii et al., 2002; Underkoffler and Ishii, 1999) Architectural Model $\times$ - - - $\times$ - - DATIBLE
@@
14. Senseboard 2001 (Jacob et al., 2002) Grid - - - $\times$ - - $\times$ CONSTNIBLE
@@
15. Illuminating Clay 2002 (Ishii et al., 2004; Piper et al., 2002) Clay Model $\times$ - - - $\times$ - - DATIBLE
@@
16. AudioPad 2002 (Patten et al., 2002, 2006) Sounds $\times$ - - - - - $\times$ DATNIBLE
@@
17. ReacTable 2003 (Jordà et al., 2007) Music $\times$ - - - - - $\times$ DATNIBLE
@@
18. IP Network Design Workbench 2003 (Kobayashi et al., 2003) Pucks - $\times$ - - - $\times$ - TOLABLE
@@
19. Query Shapes 2004 (Ichida et al., 2004) Cubes (ActiveCubes) $\times$ - - - - $\times$ - DATABLE
@@
20. TUISTER 2004 (Butz et al., 2004) TUISTER - - $\times$ - $\times$ - - OPIBLE
21. I/O Brush 2004 (Ryokai et al., 2004, 2007) Brush - $\times$ - - $\times$ - - TOLIBLE
@@
22. PICO 2005 (Patten and Ishii, 2007) Pucks (antenas) $\times$ - - - - $\times$ - DATABLE
@@
23. ArcheoTUI 2007 (Reuter et al., 2007) Props $\times$ - - - - $\times$ - DATABLE
@@
24. Slurp 2007 (Zigelbaum et al., 2008) Eyedropper - - $\times$ - $\times$ - - OPIBLE
25. GeoTUI 2008 (Couture et al., 2008) Map (cube top view) $\times$ - - - - - $\times$ DATNIBLE
@@
26. Relief 2010 (Leithinger and Ishii, 2010; Leithinger et al., 2011) 2.5D shape display (terrain) $\times$ - - - $\times$ - - DATIBLE
@@
27. Teegi 2014 (Frey et al., 2014) Teegi character (user’s activity) $\times$ - - - $\times$ - - DATIBLE
@@
28. SoundFORMS 2016 (Colter et al., 2016) Trigger pins - - $\times$ - - $\times$ - OPABLE
@@
29. reSpire 2019 (Choi et al., 2019) Shape-changing fabric $\times$ - - - $\times$ - - DATIBLE
30. CairnFORM 2019 (Daniel et al., 2019) Ring chart $\times$ - - - - $\times$ - DATABLE
31. Embodied Axes 2020 (Cordeil et al., 2020) Orthogonal arms (data axes) - - $\times$ - $\times$ - - OPIBLE
32. CoDa 2020 (Veldhuis et al., 2020) Tokens (data points) $\times$ - - - $\times$ - - DATIBLE
@@
33. SABLIER 2022 (Mahieux et al., 2022) Hourglass - - $\times$ - $\times$ - - OPIBLE
@@
(a) Relief: Terrain dynamic exploration (#26).:
(b) SABLIER: Time navigation in cultural heritage scopes (#33).:
(c) Head Prop: Neurosurgical visualization (#5).:
(d) ArcheoTUI: Virtual archaeology fragments’ assembly (#23).:
(e) GraspDraw: 2D shape edition (#6).:
(f) GeoTUI: Cutting line selection for subsoil exploration (#25).:
The hallmarks of the 33 applicative tangible user interfaces are reported in Table 7. This collection provides 28 orthogonal hallmarks. Indeed, 8 applications that share the same hallmarks are distributed into 3 clusters [^5] The three clusters’ hallmarks are and (1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0): Relief (#26), and reSpire (#29); (0, 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0): CAAD 3D Modelling System (#2) and Query Shapes (#19); (0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0): TUISTER (#20), Slurp (#24), Embodied Axes (#31), and SABLIER (#33). . The specimens in these clusters are of various tangible user interface genres. However, some specimens of the same genre (e.g., AudioPad (#16) and ReacTable (#17)) have different hallmarks, thus demonstrating that the new holistic terminology is able to discriminate physical user interfaces at the application level (i.e., beyond their morphological and technological descriptions). This collection also provides 26 orthogonal binary hallmarks (i.e., 10 applications are now distributed into the 3 clusters [^6] Pinwheels (#9) joins the cluster of Relief (#26) and reSpire (#29) because sharing the same binary hallmark: (1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0). Voodoo Dolls (#10) joins the cluster of CAAD 3D Modelling System (#2) and Query Shapes (#19) because having the same binary hallmark: (0, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0). ).
@@ 4.3. Corpus Classification
In Class III applications, the tangible or graspable part of the user interface is not about data, but rather tools or operations. For example, the I/O Brush (#21) uses World as an opible (“palette”), to draw with a tolible (“brush”) on a datnible (“canvas”). The tabletop GraspDraw (#6) edits datnibles (“shapes”) using a tolable (“pucks”) that serve as physical anchors. Finally, the tabletop GeoTUI (#25) controls a tolnible (“cutting line”) with a tolible (“ruler”) or a tolable (“two-puck prop”) on a datnible (“top view map”).
Artifacts of Class IV are designed to integrate across various applications, whose scopes can extend beyond tangible user interfaces (e.g., graphical user interfaces (Zigelbaum et al., 2008), augmented reality (Cordeil et al., 2020), and virtual reality caves (Mahieux et al., 2022)). For example, Slurp (#24) is an “eyedropper” that can extract and inject digital information. TUISTER (#20), Embodied Axes (#31), and SABLIER (#33) provide artifacts to browse hierarchical structures, control 3D positions, and navigate through space and time, respectively.
@@ 5. Mapping the Previous Terms
ASUR++ Terms (Dubois et al., 2003)
@@
Core Tangibles (Ullmer et al., 2008)
Pyfos (Shaer et al., 2004)
@@
(Shaer et al., 2004; Ullmer, 2002; Ullmer et al., 2005)
@@
Intangible representations (Ullmer, 2002; Ullmer et al., 2005)
Tangible representations (Ullmer, 2002; Ullmer et al., 2005)
Robject (Dubois et al., 2003)
@@
Passive Props (Hinckley et al., 1994)
Phicons (Ullmer and Ishii, 1997)
Artifacts (Ullmer and Ishii, 2000)
Tangibles (Ullmer and Ishii, 2000)
@@
(Shaer et al., 2004; Ullmer, 2002; Ullmer et al., 2005)
Phandles (Ullmer and Ishii, 1997)
Phidgets (Greenberg and Fitchett, 2001)
Figure 3. Mapping thirteen previous terms for physical representations and controls on the new conceptual space.:
@@ 6. Cross-Matching Classifications
Some previous classifications have discriminated between tangible user interfaces according to their morphology (i.e., how they are physically shaped, such as surfaces (Ishii, 2008), everyday objects (Ishii, 2008), or discrete tangibles (Ishii et al., 2012)) and behavior (i.e., how users interact with them and how they interact with users, such as assembly (Ishii, 2008), deformation (Ishii et al., 2012), or transformation (Ishii et al., 2012)). Figure 4 bridges these previous classifications and the tangibility classes through the collection of applicative tangible user interfaces.
Matching previous classifications, the applications’ tangibility classes are split into five clusters. The present collection of applications already shows that previous classifications and tangibility classes have no direct correspondence. Whereas “Artifacts & Objects” gathers applications coming from the four classes, “Ambient Media” and “Constructive Assemblies” count only one class each. Moreover, this initial matching may evolve when more applications are included. Indeed, even if some genres fit implicitly into certain classes well, based on their definition (e.g., interactive surfaces, such as tabletops, falign well with Class II and Class III), some classes may appear in certain clusters. For example, constructive assemblies could also be stackable audio or video filters that are tools controlling intangible data (i.e., describing such an application specimen would then require bridging Class III with “Constructive Assemblies”). Furthermore, other clusters may appear because certain genres of tangible user interfaces are absent from the collection of applications (e.g., “Tangible Telepresence” (Ishii, 2008) and “Tangibles with Kinetic Memory” (Ishii, 2008)). However, we believe that the descriptive power of the four classes may provide a sufficient design space to integrate any new applicative tangible user interface and related physical user interfaces.
@@
Symbolic tokens (Ullmer and Ishii, 2000)
@@
Transformable continuous tangibles (Ishii et al., 2012)
@@
Deformable continuous tangibles (Ishii et al., 2012)
@@
(Ishii, 2008)
(Ishii, 2008)
(Ishii, 2008)
(Ishii, 2008)
(Ishii, 2008)
(Ishii, 2008)
Figure 4. Matching previous classifications with what–how tangibility classes through a collection of applications.:
@@ 7. Limitations and Future Work
A limitation of this work is that analyses were not consolidated with the specimens’ designers. Describing the 33 applications was mostly straightforward, but sometimes required advanced questioning. An example is when describing the Pinwheels specimen (#9), which represents information flow in an ambient manner. At first glance, Pinwheels must be datnible: airflow represents the datum flow. However, this idea evolved from using airflow in the ambientROOM (Ishii et al., 1998)—merely needing air coming into the room through a hole—to using visual “spinning pinwheels” (Dahley et al., 1998; Ishii et al., 2001b; Wisneski et al., 1998). Hence, Pinwheels are described as datibles [^7] Treating single datnibles (airflow) as physical user interfaces would lead to reconsidering Class I definition or to creating a new class.. Another example lies in the vertical measurement graphs (i.e., usage rate, response time, and running cost graphs) placed next to the IP Network Design Workbench (#18). Some could have considered these graphs as auxiliary GUIs. However, they were considered as opnibles (i.e., intangible measurement operations), as well as measurements in the Urp (#13) that were considered as operations. Advanced questioning can become tricky when design thoughts are not clearly detailed in the articles. The community must benefit from future work that describes previous physical user interface specimens in a publicly shared database (e.g., (Fleck et al., 2018)) and future articles that describe specimens using holistic terminologies.
@@
This work focused on the roles found in tangible and related physical user interfaces without considering remote collaboration conditions (Brave et al., 1998). However, in such conditions, remote users may be represented by the user interface through streams (Ishii, 1990; Ishii and Kobayashi, 1992; Leithinger et al., 2014), shapes (Brave and Dahley, 1997; Brewer et al., 2007), or avatars (Lee et al., 2025) and even interact with physical entities (Brave and Dahley, 1997; Leithinger et al., 2014). Therefore, these conditions raise the question of defining a role for remote users specifically (or considering them merely as data or tools?), refining their body parts (e.g., head, forearms, and hands), and defining terms that describe how they appear in the user interface. These questions may extend the terminology to interactions with robots and cobots in Human-Robot Interaction (Stoeva et al., 2024) and Extended Reality conditions (Gonzalez-Sanchez and Guerra-Silva, 2025).
@@
The description of bodied entities is refined on only two levels of embodiment (i.e., tangible and graspable), which roughly simplifies the twenty values of Fishkin’ taxonomy (Fishkin, 2004). However, this article considered a level of embodiment for bodiless entities (i.e., intangible), even if intangible entities can also encounter various levels of representativeness (e.g., a beep or slang audio feedback when sending a file to trash, versus paper-creasing noise).
@@
This work separated hallmarks into classes. However, defining a distance calculation between hallmarks could serve the needs of automatic classification algorithms (e.g., cladistics (Fleck et al., 2018)) that would help sort, organize, and browse the vast collection of specimens provided by the field of tangible user interfaces over the past three decades.
@@
This article introduces the concept of interactional entities briefly, which future work will need to define more extensively.

### v3

@@ 10. Voodoo Dolls 1998 (isidoro1998voodoo) Spongy object $\times$ - - - - $\times$ - DATABLE
11. mediaBlocks 1998 (ullmer1999mediablocks; ullmer1998mediablocks) mediaBlock - - $\times$ - - $\times$ - OPABLE
@@
12. musicBottles 1999 (ishii1999musicbottles; ishii2001musicbottles) Music $\times$ - - - - - $\times$ DATNIBLE
@@
13. Urp (Urban Planning Workbench) 1999 (benjoseph2001urp; ishii2002urp; underkoffler1999urp) Architectural Model $\times$ - - - $\times$ - - DATIBLE
@@
14. Senseboard 2001 (jacob2002senseboard) Grid - - - $\times$ - - $\times$ CONSTNIBLE
@@
15. Illuminating Clay 2002 (ishii2004illclay; piper2002illclay) Clay Model $\times$ - - - $\times$ - - DATIBLE
@@
16. AudioPad 2002 (patten2002audiopad; patten2006audiopad) Sounds $\times$ - - - - - $\times$ DATNIBLE
@@
17. ReacTable 2003 (jorda2007reactable) Music $\times$ - - - - - $\times$ DATNIBLE
@@
18. IP Network Design Workbench 2003 (kobayashi2003ip) Pucks - $\times$ - - - $\times$ - TOLABLE
@@
19. Query Shapes 2004 (ichida2004retrieval) Cubes (ActiveCubes) $\times$ - - - - $\times$ - DATABLE
@@
20. TUISTER 2004 (butz2004tuister) TUISTER - - $\times$ - $\times$ - - OPIBLE
21. I/O Brush 2004 (ryokai2004iobrush; ryokai2007iobrush) Brush - $\times$ - - $\times$ - - TOLIBLE
@@
22. PICO 2005 (patten2007pico) Pucks (antenas) $\times$ - - - - $\times$ - DATABLE
@@
23. ArcheoTUI 2007 (reuter2007archeotui) Props $\times$ - - - - $\times$ - DATABLE
@@
24. GeoTUI 2008 (couture2008geotui) Map (cube top view) $\times$ - - - - - $\times$ DATNIBLE
@@
25. Slurp 2008 (zigelbaum2008slurp) Eyedropper - - $\times$ - $\times$ - - OPIBLE
26. Relief 2010 (leithinger2010relief; leithinger2011relief) 2.5D shape display (terrain) $\times$ - - - $\times$ - - DATIBLE
@@ 25. Slurp 2008 (zigelbaum2008slurp) Eyedropper - - $\times$ - $\times$ - - OPIBLE
27. Teegi 2014 (frey2014teegi) Teegi character (user’s activity) $\times$ - - - $\times$ - - DATIBLE
@@
28. SoundFORMS 2016 (colter2016soundforms) Trigger pins - - $\times$ - - $\times$ - OPABLE
@@
29. CairnFORM 2019 (daniel2019cairnform) Ring chart $\times$ - - - - $\times$ - DATABLE
30. reSpire 2019 (choi2019respire) Shape-changing fabric $\times$ - - - $\times$ - - DATIBLE
31. Embodied Axes 2020 (cordeil2020axes) Orthogonal arms (data axes) - - $\times$ - $\times$ - - OPIBLE
32. CoDa 2020 (veldhuis2020coda) Tokens (data points) $\times$ - - - $\times$ - - DATIBLE
@@
33. SABLIER 2022 (mahieux2022sablier) Hourglass - - $\times$ - $\times$ - - OPIBLE
@@
The hallmarks of the 33 applicative tangible user interfaces are reported in Table 7. This collection provides 28 orthogonal hallmarks. Indeed, 8 applications that share the same hallmarks are distributed into 3 clusters [^5] The three clusters’ hallmarks are and (1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0): Relief (#26), and reSpire (#29); (0, 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0): CAAD 3D Modelling System (#2) and Query Shapes (#19); (0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0): TUISTER (#20), Slurp (#25), Embodied Axes (#31), and SABLIER (#33). . The specimens in these clusters are of various tangible user interface genres. However, some specimens of the same genre (e.g., AudioPad (#16) and ReacTable (#17)) have different hallmarks, thus demonstrating that the new holistic terminology is able to discriminate physical user interfaces at the application level (i.e., beyond their morphological and technological descriptions). This collection also provides 26 orthogonal binary hallmarks (i.e., 10 applications are now distributed into the 3 clusters [^6] Pinwheels (#9) joins the cluster of Relief (#26) and reSpire (#29) because sharing the same binary hallmark: (1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0). Voodoo Dolls (#10) joins the cluster of CAAD 3D Modelling System (#2) and Query Shapes (#19) because having the same binary hallmark: (0, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0). ).
@@ 4.3. Corpus Classification
In Class III applications, the tangible or graspable part of the user interface is not about data, but rather tools or operations. For example, the I/O Brush (#21) uses World as an opible (“palette”), to draw with a tolible (“brush”) on a datnible (“canvas”). The tabletop GraspDraw (#6) edits datnibles (“shapes”) using a tolable (“pucks”) that serve as physical anchors. Finally, the tabletop GeoTUI (#24) controls a tolnible (“cutting line”) with a tolible (“ruler”) or a tolable (“two-puck prop”) on a datnible (“top view map”).
Artifacts of Class IV are designed to integrate across various applications, whose scopes can extend beyond tangible user interfaces (e.g., graphical user interfaces (zigelbaum2008slurp), augmented reality (cordeil2020axes), and virtual reality caves (mahieux2022sablier)). For example, Slurp (#25) is an “eyedropper” that can extract and inject digital information. TUISTER (#20), Embodied Axes (#31), and SABLIER (#33) provide artifacts to browse hierarchical structures, control 3D positions, and navigate through space and time, respectively.
@@ 5. Mapping the Previous Terms
ASUR++ Terms (dubois2003asur++)
@@
Core Tangibles (ullmer2008core)
Pyfos (shaer2004tac)
@@
(shaer2004tac; ullmer2002phdthesis; ullmer2005token)
@@
Intangible representations (ullmer2002phdthesis; ullmer2005token)
Tangible representations (ullmer2002phdthesis; ullmer2005token)
Robject (dubois2003asur++)
@@
Passive Props (hinckley1994passiveprops)
Phicons (ullmer1997metadesk)
Artifacts (ullmer2000emerging)
Tangibles (ullmer2000emerging)
@@
(shaer2004tac; ullmer2002phdthesis; ullmer2005token)
Phandles (ullmer1997metadesk)
Phidgets (greenberg2001phidgets)
Figure 3. Mapping thirteen previous terms for physical representations and controls on the new conceptual space.
The infography is a graphical projection of the previous terms on a two-dimensional space determined by the what–how taxonomy that originated the terms of the new terminology. The horizontal axis is composed of the four values about what digital role is assigned: Datum, Tool, Operation, or Constraint. The vertical axis is composed of three tangibility degrees: Tangible, Graspable, and Intangible. The resulting grid is covered by rectangles according to previous terms mapping on the two axes.
@@ 6. Cross-Matching Classifications
Some previous classifications have discriminated between tangible user interfaces according to their morphology (i.e., how they are physically shaped, such as surfaces (ishii2008beyond), everyday objects (ishii2008beyond), or discrete tangibles (ishii2012radical)) and behavior (i.e., how users interact with them and how they interact with users, such as assembly (ishii2008beyond), deformation (ishii2012radical), or transformation (ishii2012radical)). Figure 4 bridges these previous classifications and the tangibility classes through the collection of applicative tangible user interfaces.
@@
Symbolic tokens (ullmer2000emerging)
@@
Transformable continuous tangibles (ishii2012radical)
@@
Deformable continuous tangibles (ishii2012radical)
@@
Figure 4. Matching previous classifications with what–how tangibility classes through a collection of applications.
The diagram is composed of four columns: (1) Genre, (2) Subgenre, (3) Application, and (4) Tangibility Class. From left to right, five genres of applications are linked to thirteen subgenres, which are linked to the 33 applications. Then, the application are linked to their corresponding classes.
Matching previous classifications, the applications’ tangibility classes are split into five clusters. The present collection of applications already shows that previous classifications and tangibility classes have no direct correspondence. Whereas “Artifacts & Objects” gathers applications coming from the four classes, “Ambient Media” and “Constructive Assemblies” count only one class each. Moreover, this initial matching may evolve when more applications are included. Indeed, even if some genres fit implicitly into certain classes well, based on their definition (e.g., interactive surfaces, such as tabletops, falign well with Class II and Class III), some classes may appear in certain clusters. For example, constructive assemblies could also be stackable audio or video filters that are tools controlling intangible data (i.e., describing such an application specimen would then require bridging Class III with “Constructive Assemblies”). Furthermore, other clusters may appear because certain genres of tangible user interfaces are absent from the collection of applications (e.g., “Tangible Telepresence” (ishii2008beyond) and “Tangibles with Kinetic Memory” (ishii2008beyond)). However, we believe that the descriptive power of the four classes may provide a sufficient design space to integrate any new applicative tangible user interface and related physical user interfaces.
@@ 7. Limitations and Future Work
A limitation of this work is that analyses were not consolidated with the specimens’ designers. Describing the 33 applications was mostly straightforward, but sometimes required advanced questioning. An example is when describing the Pinwheels specimen (#9), which represents information flow in an ambient manner. At first glance, Pinwheels must be datnible: airflow represents the datum flow. However, this idea evolved from using airflow in the ambientROOM (ishii1998ambientroom)—merely needing air coming into the room through a hole—to using visual “spinning pinwheels” (dahley1998ambient; ishii2001pinwheels; wisneski1998pinwheels). Hence, Pinwheels are described as datibles [^7] Treating single datnibles (airflow) as physical user interfaces would lead to reconsidering Class I definition or to creating a new class.. Another example lies in the vertical measurement graphs (i.e., usage rate, response time, and running cost graphs) placed next to the IP Network Design Workbench (#18). Some could have considered these graphs as auxiliary GUIs. However, they were considered as opnibles (i.e., intangible measurement operations), as well as measurements in the Urp (#13) that were considered as operations. Advanced questioning can become tricky when design thoughts are not clearly detailed in the articles. The community must benefit from future work that describes previous physical user interface specimens in a publicly shared database (e.g., (fleck2018cladistics)) and future articles that describe specimens using holistic terminologies.
@@
This work focused on the roles found in tangible and related physical user interfaces without considering remote collaboration conditions (brave1998remote). However, in such conditions, remote users may be represented by the user interface through streams (ishii1990teamworkstation; ishii1992clearboard; leithinger2014inform), shapes (brave1997intouch; brewer2007nimio), or avatars (lee2025since) and even interact with physical entities (brave1997intouch; leithinger2014inform). Therefore, these conditions raise the question of defining a role for remote users specifically (or considering them merely as data or tools?), refining their body parts (e.g., head, forearms, and hands), and defining terms that describe how they appear in the user interface. These questions may extend the terminology to interactions with robots and cobots in Human-Robot Interaction (stoeva2024body) and Extended Reality conditions (gonzalez2025exploring).
@@
The description of bodied entities is refined on only two levels of embodiment (i.e., tangible and graspable), which roughly simplifies the twenty values of Fishkin’ taxonomy (fishkin2004taxonomy). However, this article considered a level of embodiment for bodiless entities (i.e., intangible), even if intangible entities can also encounter various levels of representativeness (e.g., a beep or slang audio feedback when sending a file to trash, versus paper-creasing noise).
@@
This work separated hallmarks into classes. However, defining a distance calculation between hallmarks could serve the needs of automatic classification algorithms (e.g., cladistics (fleck2018cladistics)) that would help sort, organize, and browse the vast collection of specimens provided by the field of tangible user interfaces over the past three decades.
@@
Finally, this article introduces the concept of interactional entities briefly, which future work will need to define more extensively.